OpenStudy (anonymous):

The length and width of a rectangle are in the ratio of 3:7. If the area of the rectangle is 525m^2, find its dimensions. (Explain how you got the answer.)

OpenStudy (anonymous):

3:7 we can write x there and multiply width and length \[3x\cdot7x=525\]

OpenStudy (anonymous):

oh--Thankyou............so it would be 21x^2 = 525 and solve for x..right?

OpenStudy (anonymous):

Solve the following for L and w:\[\left\{\frac{3}{7}=\frac{L}{w},L w=525\right\} \]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yeah and when you find x multiply it by 5 for length and by 7 for width or do like @robtobey answered

OpenStudy (anonymous):

i still dont get it

OpenStudy (anonymous):

can u please explain it again?

OpenStudy (anonymous):

how do we know which one is length or width if we just use x? can you please explain it again...cuz i dont get it.